the sum of the ages of joe and jim is 77. eight years from now, joe will be twice jim's age. find joe's present age?
Answer by shree840(260)

x=joe
y=jim
x+y=77
x+8=2(y+8)or
x-2y=8
------------
subtract line 2 fom line 1
x+y=77
x-2y=8 subtracting the new line is
-x+2y=-8
y+2y=77-8=69
3y=69
y=23
x=77-23=54
answer joe=54 jim=31